OpenStudy (anonymous):

will medalllll help The function f(x) = 467(5)x represents the growth of a ladybug population every year in a wooded area. Adrianne wants to manipulate the formula to an equivalent form that calculates every 3 months, not every year. Which function is correct for Adrianne's purposes? f(x) = 67(5)^x f(x) = 467(5^12)^x/12 f(x) = 467(5^1/4)^4x f(x) = 4672(5)^x

OpenStudy (anonymous):

\[f(x)=67(5)^{x}\] \[f(x)=467(5^{12})^{x/12}\] \[f(x)=467(5^{1/4})^{4x}\] \[f(x)=4672(5)^{x}\]

OpenStudy (bibby):

\( f(x) = 67(5)^x \\ f(x) = 467(5^{12})^{\frac{x}{12}} \\ f(x) = 467(5^{\frac{1}{4}})^{4x} \\ f(x) = 4672(5)^x \) is that it? oh you rewrote it yourself

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

are those possible answer choices?

OpenStudy (anonymous):

Yes they are

OpenStudy (misty1212):

that is what i thought

OpenStudy (misty1212):

every three months is one quarter of a year

OpenStudy (anonymous):

So then it would be C?

OpenStudy (misty1212):

the number in front does not change

OpenStudy (misty1212):

yeah as always it is C

OpenStudy (anonymous):

Thank you :)